# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

### v2

@@ 6.5. Detecting Klein bottles and Dehn surgeries on $\mathcal{L}$ and $L_{B}$
(3) For each hyperbolic vertex $v\in V(\Gamma_{M})$, $f_{v}:\widehat{\pi_{1}M_{v}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{f^{\dagger}(v)}}$ respects the peripheral structure according to [57, Lemma 6.3] (or Theorem 3.2 (1)). In fact, according to the commutative diagram in Proposition A.3 (3), $f_{v}$ matches up the boundary components of $M_{v}$ coming from the JSJ-tori (i.e. corresponding to the edges $e\in E(\Gamma_{M})$ adjoining $v$) to the boundary components of $N_{f^{\dagger}(v)}$ coming from the JSJ-tori (i.e. corresponding to the edges $f^{\dagger}(e)$). Thus, $f_{v}$ matches up the remaining boundary components, i.e. matches up $M_{v}\cap\partial M$ with $N_{f^{\dagger}(v)}\cap\partial N$. Therefore, up to reordering, $f_{v}$ sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N_{f^{\dagger}(v)}}$. The commutative diagram in Proposition A.3 (3) then implies that $f$ in fact sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N}$.
(4) Based on our proof for (3), (4) follows from the fact that $f_{v}$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at $\partial_{i}^{hyp}M$ (Theorem 3.2 (2)), together with the commutative diagram Proposition A.3 (3). ∎

### v3

@@ 6.5. Detecting Klein bottles and Dehn surgeries on $\mathcal{L}$ and $L_{B}$
(3) For each hyperbolic vertex $v\in V(\Gamma_{M})$, $f_{v}:\widehat{\pi_{1}M_{v}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{f^{\dagger}(v)}}$ respects the peripheral structure according to [56, Lemma 6.3]. In fact, according to Proposition A.3 (4), $f_{v}$ matches up the boundary components of $M_{v}$ coming from the JSJ-tori (i.e. corresponding to the edges $e\in E(\Gamma_{M})$ adjoining $v$) to the boundary components of $N_{f^{\dagger}(v)}$ coming from the JSJ-tori (i.e. corresponding to the edges $f^{\dagger}(e)$). Thus, $f_{v}$ matches up the remaining boundary components, i.e. matches up $M_{v}\cap\partial M$ with $N_{f^{\dagger}(v)}\cap\partial N$. Therefore, up to reordering, $f_{v}$ sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N_{f^{\dagger}(v)}}$. The commutative diagram in Proposition A.3 (3) then implies that $f$ in fact sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N}$.
(4) Based on our proof for (3), item (4) now follows from the fact that $f_{v}$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at $\partial_{i}^{hyp}M$ (Theorem 3.2), together with the commutative diagram in Proposition A.3 (3). ∎